# NeuralFuse: Learning to Recover the Accuracy of Access-Limited Neural Network Inference in Low-Voltage Regimes

### v2

@@ F.2 GTSRB
Base Model NF C.A. 1% B.E.R. 0.5% B.E.R. P.A. C.A. (NF) P.A. (NF) R.P. P.A. C.A. (NF) P.A. (NF) R.P. ResNet18 ConvL 95.5 95.7 91.1 4.7 54.2 93.4 89.5 1.9 14.3 ConvS 94.4 68.6 12 31.7 94.8 87.7 4.2 12.4 DeConvL 36.9 95.6 91.3 4.3 54.4 75.2 95.4 93.4 1.1 18.1 DeConvS 16.0 95.7 78.1 9.1 41.2 12.7 95.8 90.1 3.3 14.9 UNetL 96.2 93.8 1.0 56.9 96.2 93.5 1.6 18.3 UNetS 95.9 85.1 6.9 48.2 95.5 91.4 2.8 16.2 ResNet50 ConvL 95.0 95.6 71.6 20 42.1 94.6 90.6 3.7 16.6 ConvS 94.8 50.5 22 21.0 95.4 84.5 8.5 10.5 DeConvL 29.5 94.9 71.6 21 42.0 74.0 94.7 91.6 2.9 17.6 DeConvS 16.9 93.0 56.4 17 26.9 13.0 94.6 87.4 5.9 13.5 UNetL 94.5 80.6 15 51.1 96.5 93.7 2.3 19.7 UNetS 94.7 64.7 22 35.2 95.9 90.6 4.8 16.7 VGG11 ConvL 91.9 94.8 85.7 7.2 50.9 93.9 92.6 0.7 27.7 ConvS 91.1 62.2 11 27.3 90.9 80.5 3.5 15.7 DeConvL 34.9 95.0 84.6 7.6 49.7 64.9 93.6 91.9 0.6 27.1 DeConvS 12.4 92.4 67.5 11 32.6 10.8 92.3 83.1 3.7 18.2 UNetL 92.2 83.2 6.0 48.3 94.8 90.6 1.7 25.7 UNetS 94.7 73.4 10 38.5 94.6 88.9 2.2 24.1 VGG16 ConvL 95.2 96.3 72.4 12 57.3 95.6 93.2 1.8 34.4 ConvS 94.1 39.8 13 24.6 94.3 82.2 6.2 23.4 DeConvL 15.1 96.4 72.0 12 56.9 58.8 95.6 93.1 2.0 34.3 DeConvS 6.8 93.8 50.9 13 35.8 8.9 95.1 84.0 5.3 25.2 UNetL 95.8 78.6 11 63.5 96.0 92.8 2.0 34.0 UNetS 94.3 63.3 14 48.1 95.4 87.8 3.6 29.0 VGG19 ConvL 95.5 96.0 88.3 7.2 51.7 95.6 93.4 2.1 24.2 ConvS 93.8 69.0 14 32.4 94.9 87.0 4.4 17.8 DeConvL 36.6 95.4 87.2 7.5 50.6 69.1 95.5 92.4 2.2 23.3 DeConvS 6.8 94.5 73.1 12 36.5 11.1 95.5 88.8 3.7 19.7 UNetL 95.4 88.2 6.7 51.7 94.9 91.7 2.5 22.6 UNetS 94.6 80.6 9.0 44.1 96.5 90.8 3.4 21.6 • [Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
@@ F.3 ImageNet-10
[Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
@@ F.4 CIFAR-100
Base Model NF C.A. 1% B.E.R. 0.5% B.E.R. 0.35% B.E.R. P.A. C.A. (NF) P.A. (NF) R.P. P.A. C.A. (NF) P.A. (NF) R.P. P.A. C.A. (NF) P.A. (NF) R.P. ResNet18 ConvL 73.7 54.8 11.0 7.7 6.4 65.2 39.0 7.1 18.1 69.4 42.9 6.2 11.4 ConvS 49.7 4.2 2.2 -0.4 70.0 24.5 7.6 3.6 72.1 35.1 7.3 3.7 DeConvL 4.6 55.2 11.9 8.2 7.3 20.9 66.3 38.2 6.9 17.3 31.4 69.2 42.9 5.5 11.4 DeConvS 2.9 32.7 4.0 2.2 -0.6 7.4 68.2 25.9 6.8 5 7.6 71.6 35.8 5.5 4.4 UNetL 50.6 14.5 8.9 10.0 66.2 40.1 6.4 19.2 70.3 46.3 5.5 14.9 UNetS 26.8 4.6 2.5 -0.0 67.1 28.8 6.8 7.9 70.9 38.3 6.4 6.9 ResNet50 ConvL 73.5 63.5 3.2 1.7 0.1 68.4 28.8 6.7 7.6 72.0 40.8 7.5 5.1 ConvS 65.5 3.2 1.6 0.1 71.9 23.1 6.9 1.9 73.0 37.4 8.0 1.7 DeConvL 3.0 59.6 3.2 1.7 0.2 21.3 68.1 28.6 7.0 7.4 35.7 71.7 41.7 7.7 6.1 DeConvS 1.8 61.1 3.2 1.7 0.1 7.0 70.3 25.0 6.7 3.7 8.6 72.8 38.9 7.9 3.3 UNetL 39.0 5.0 1.7 1.9 66.6 36.5 6.2 15.3 70.8 45.3 6.7 9.6 UNetS 47.7 3.4 1.8 0.3 69.1 26.1 6.6 4.8 72.6 39.6 7.8 3.9 VGG11 ConvL 64.8 58.3 19.7 11 11.5 63.1 38.8 9.3 15.0 63.9 42.4 9.0 11.1 ConvS 56.6 10.4 7.4 2.2 62.7 27.9 10 4.0 63.9 41.8 8.3 10.5 DeConvL 8.2 60.3 21.2 11 13.0 23.9 63.9 40.0 9.0 16.2 31.3 64.0 42.8 9.1 11.5 DeConvS 5.7 58.3 11.8 7.9 3.5 9.4 61.9 29.8 9.9 5.9 10 63.5 36.1 10 4.8 UNetL 51.1 22.1 8.2 13.9 61.8 37.8 9.0 13.9 63.5 40.9 9.3 9.6 UNetS 51.9 13.1 7.9 4.9 61.7 29.8 9.7 6.0 63.8 35.7 9.9 4.5 VGG16 ConvL 67.8 51.4 19.2 6.0 12.6 61.8 41.1 5.6 18.7 64.9 44.9 5.3 13.8 ConvS 44.3 6.7 2.3 0.1 63.8 27.5 6.8 5.1 66.0 36.3 6.1 5.1 DeConvL 7.0 53.1 20.8 6.2 14.2 22.4 62.8 42.1 5.5 19.8 31.1 65.0 46.6 5.2 15.5 DeConvS 3.5 23.5 4.8 1.7 -1.8 7.0 62.1 29.9 6.7 7.5 7.2 64.9 38.1 6.3 7.0 UNetL 50.2 25.3 1.7 18.7 61.7 41.3 5.0 18.9 64.8 46.8 4.6 15.7 UNetS 27.7 9.9 2.1 3.3 61.6 31.3 6.3 8.9 65.0 39.8 5.9 8.7 VGG19 ConvL 67.8 59.4 29.2 8.1 18.6 65.6 46.5 6.8 12.5 66.9 49.2 7.4 7.0 ConvS 63.7 14.4 5.1 3.8 66.6 38.3 6.8 4.2 67.7 45.3 8.5 3.2 DeConvL 10.6 60.1 29.6 8.5 19.0 34.0 65.7 46.9 7.1 12.9 42.1 67.3 49.8 7.6 7.6 DeConvS 4.3 60.9 16.1 6.0 5.6 9.6 66.5 39.0 3.7 5.0 9.4 67.7 45.7 8.4 3.6 UNetL 58.7 30.2 8.2 19.6 65.5 46.9 6.5 12.9 67.4 50.0 7.5 7.9 UNetS 59.1 18.0 6.2 7.4 66.3 40.1 8.0 6.1 67.5 46.6 8.4 4.5 • [Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
@@ Appendix G Additional Experimental Results on Restricted Access Settings (Transferability)
We find that using VGG19 as a white-box surrogate model has better transferability than ResNet18 for all datasets. In addition, we can observe that some NeuralFuse generators have downward applicability if base models have a similar architecture. In other words, if we try to transfer a generator trained on a large B.E.R. (e.g., 1%) to a model with a small B.E.R. (e.g., 0.5%), the performance will be better than that of a generator trained with the original B.E.R. (e.g., 0.5%). For example, in Table 16, we could find that if we use VGG19 as a source model to train the generator ConvL (1%), the generator could deliver better performance (in terms of P.A. (NF)) when applied to similar base models (e.g., VGG11, VGG16, or VGG19) under a 0.5% B.E.R., compared to using itself as a source model (shown in Table 12). We conjecture that this is because the generators trained on a larger B.E.R. can also cover the error patterns of a smaller B.E.R., and thus they have better generalizability across smaller B.E.Rs.
@@ G.1 CIFAR-10
The results of CIFAR-10 in which NeuralFuse is trained at 1% B.E.R. are shown in Table 16.
S.M. T.M. B.E.R. C.A. P.A. ConvL (1%) UNetL (1%) C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 ResNet18 0.5% 92.6 70.1 11.6 89.8 89.5 0.2 19.4 86.6 86.2 0.3 16.1 ResNet50 1% 92.6 26.1 9.4 89.5 36.0 19 9.9 85.2 38.8 19 12.7 0.5% 61.0 10.3 75.1 10 14.1 77.1 5.0 16.1 VGG11 1% 88.4 42.2 11.6 88.4 62.5 8.4 20.3 76.8 61.1 8.5 18.9 0.5% 63.6 9.3 81.0 4.6 17.4 73.7 3.0 10.1 VGG16 1% 90.3 35.7 7.9 89.6 63.3 18 27.6 85.2 59.9 16 24.2 0.5% 66.6 8.1 85.0 3.4 18.4 80.2 4.5 13.6 VGG19 1% 90.5 36.0 12.0 89.6 50.7 22 14.7 85.3 51.1 16 15.1 0.5% 64.2 12.4 80.2 8.7 16.0 76.5 7.8 12.3 VGG19 ResNet18 1% 92.6 38.9 12.4 89.8 61.0 17 22.1 87.0 69.7 11 30.8 0.5% 70.1 11.6 86.1 6.9 16.0 84.2 3.0 14.1 ResNet50 1% 92.6 26.1 9.4 89.9 34.0 19 7.9 87.0 44.2 17 18.1 0.5% 61.0 10.3 76.5 10 15.5 80.7 4.2 19.7 VGG11 1% 88.4 42.2 11.6 89.7 76.5 7.0 34.3 87.1 79.9 5.6 37.7 0.5% 63.6 9.3 88.0 2.1 24.4 85.4 0.8 21.8 VGG16 1% 90.3 35.7 7.9 89.6 75.5 12 39.8 87.2 78.9 7.8 43.2 0.5% 66.6 8.1 88.9 0.6 22.3 86.2 0.3 19.6 VGG19 0.5% 90.5 64.2 12.4 89.8 89.6 8.7 25.4 87.4 86.8 0.4 22.6 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
@@ G.2 GTSRB
In Tables 17 and 18, we show the results on GTSRB in which NeuralFuse is trained at 1.5% and 1% B.E.R., respectively.
S.M. T.M. B.E.R. C.A. P.A. ConvL (1.5%) UNetL (1.5%) C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 ResNet18 1% 95.5 36.9 16.0 95.7 93.9 1.9 57.0 94.9 94.4 0.4 57.5 0.5% 75.2 12.7 95.7 0.2 20.5 94.8 0.2 19.6 ResNet50 1% 95.0 29.5 16.9 94.4 37.0 22 7.5 94.4 47.1 23 17.6 0.5% 74.0 13.0 77.5 13 3.5 84.8 9.5 10.8 VGG11 1% 91.9 34.9 12.4 92.8 45.2 10 10.3 91.4 50.5 13 15.6 0.5% 64.9 10.8 79.4 5.8 14.5 83.9 4.2 19.0 VGG16 1% 95.2 15.1 6.8 95.4 31.1 13 15.9 94.6 36.8 12 21.7 0.5% 58.8 8.9 84.5 8.3 25.8 86.0 8.6 27.2 VGG19 1% 95.5 36.6 6.8 95.0 56.4 15 19.8 94.3 60.8 15 24.2 0.5% 69.1 11.1 86.9 3.4 17.8 87.7 3.8 18.6 VGG19 ResNet18 1% 95.5 36.9 16.0 88.4 50.3 12 13.4 92.8 63.7 16 26.8 0.5% 75.2 12.7 77.9 7.4 2.7 87.5 3.9 12.3 ResNet50 1% 95.0 29.5 16.9 87.5 29.7 17 0.2 92.5 40.4 21 10.9 0.5% 74.0 13.0 67.9 17 -6.1 77.5 15 3.5 VGG11 1% 91.9 34.9 12.4 89.7 47.1 11 12.2 93.5 60.0 12 25.1 0.5% 64.9 10.8 76.3 5.1 11.4 86.0 3.8 21.1 VGG16 1% 95.2 15.1 6.8 93.0 29.2 15 14.1 93.0 38.5 16 23.4 0.5% 58.8 8.9 75.7 12 16.9 79.9 8.3 21.1 VGG19 1% 95.5 36.6 6.8 95.1 87.4 6.0 50.8 94.6 88.7 5.0 52.1 0.5% 69.1 11.1 92.4 2.4 23.3 92.4 2.2 23.3 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
S.M. T.M. B.E.R. C.A. P.A. ConvL (1%) UNetL (1%) C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 ResNet18 0.5% 95.5 75.2 12.7 95.7 95.3 0.5 20.1 96.2 95.7 0.3 20.5 ResNet50 1% 95.0 29.5 16.9 94.5 35.6 21 6.1 95.6 42.6 23 13.1 0.5% 74.0 13.0 78.8 13 4.8 87.3 9.0 13.3 VGG11 1% 91.9 34.9 12.4 93.1 45.8 11 10.9 94.0 47.1 14 12.2 0.5% 64.9 10.8 81.8 5.0 16.9 84.2 4.8 19.3 VGG16 1% 95.2 15.1 6.8 95.5 26.5 12 11.4 95.5 32.4 11 17.3 0.5% 58.8 8.9 82.2 9.0 23.4 85.4 6.7 26.6 VGG19 1% 95.5 36.6 6.8 94.9 53.2 14 16.6 95.6 60.9 15 24.3 0.5% 69.1 11.1 85.4 4.5 16.3 87.5 3.7 18.4 VGG19 ResNet18 1% 95.5 36.9 16.0 93.7 53.1 16 16.2 95.0 63.4 18 26.5 0.5% 75.2 12.7 83.9 7.6 8.7 89.7 4.8 14.5 ResNet50 1% 95.0 29.5 16.9 92.8 30.6 18 1.1 95.4 38.9 22 9.4 0.5% 74.0 13.0 74.7 18 0.7 81.5 16 7.5 VGG11 1% 91.9 34.9 12.4 93.7 50.6 11 15.7 95.1 58.9 15 24.0 0.5% 64.9 10.8 82.3 5.1 17.4 87.5 3.7 22.6 VGG16 1% 95.2 15.1 6.8 95.2 27.8 15 12.7 95.2 33.5 14 18.4 0.5% 58.8 8.9 79.0 12 20.2 81.8 7.8 23.0 VGG19 0.5% 95.5 69.1 11.1 96.0 94.0 2.2 24.9 95.4 93.9 2.1 24.8 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.

### v3

@@ F.2 GTSRB
Base Model NF CA 1% BER 0.5% BER PA CA (NF) PA (NF) RP PA CA (NF) PA (NF) RP ResNet18 ConvL 95.5 95.7 91.1 4.7 54.2 93.4 89.5 1.9 14.3 ConvS 94.4 68.6 12 31.7 94.8 87.7 4.2 12.4 DeConvL 36.9 95.6 91.3 4.3 54.4 75.2 95.4 93.4 1.1 18.1 DeConvS 16.0 95.7 78.1 9.1 41.2 12.7 95.8 90.1 3.3 14.9 UNetL 96.2 93.8 1.0 56.9 96.2 93.5 1.6 18.3 UNetS 95.9 85.1 6.9 48.2 95.5 91.4 2.8 16.2 ResNet50 ConvL 95.0 95.6 71.6 20 42.1 94.6 90.6 3.7 16.6 ConvS 94.8 50.5 22 21.0 95.4 84.5 8.5 10.5 DeConvL 29.5 94.9 71.6 21 42.0 74.0 94.7 91.6 2.9 17.6 DeConvS 16.9 93.0 56.4 17 26.9 13.0 94.6 87.4 5.9 13.5 UNetL 94.5 80.6 15 51.1 96.5 93.7 2.3 19.7 UNetS 94.7 64.7 22 35.2 95.9 90.6 4.8 16.7 VGG11 ConvL 91.9 94.8 85.7 7.2 50.9 93.9 92.6 0.7 27.7 ConvS 91.1 62.2 11 27.3 90.9 80.5 3.5 15.7 DeConvL 34.9 95.0 84.6 7.6 49.7 64.9 93.6 91.9 0.6 27.1 DeConvS 12.4 92.4 67.5 11 32.6 10.8 92.3 83.1 3.7 18.2 UNetL 92.2 83.2 6.0 48.3 94.8 90.6 1.7 25.7 UNetS 94.7 73.4 10 38.5 94.6 88.9 2.2 24.1 VGG16 ConvL 95.2 96.3 72.4 12 57.3 95.6 93.2 1.8 34.4 ConvS 94.1 39.8 13 24.6 94.3 82.2 6.2 23.4 DeConvL 15.1 96.4 72.0 12 56.9 58.8 95.6 93.1 2.0 34.3 DeConvS 6.8 93.8 50.9 13 35.8 8.9 95.1 84.0 5.3 25.2 UNetL 95.8 78.6 11 63.5 96.0 92.8 2.0 34.0 UNetS 94.3 63.3 14 48.1 95.4 87.8 3.6 29.0 VGG19 ConvL 95.5 96.0 88.3 7.2 51.7 95.6 93.4 2.1 24.2 ConvS 93.8 69.0 14 32.4 94.9 87.0 4.4 17.8 DeConvL 36.6 95.4 87.2 7.5 50.6 69.1 95.5 92.4 2.2 23.3 DeConvS 6.8 94.5 73.1 12 36.5 11.1 95.5 88.8 3.7 19.7 UNetL 95.4 88.2 6.7 51.7 94.9 91.7 2.5 22.6 UNetS 94.6 80.6 9.0 44.1 96.5 90.8 3.4 21.6 • Note. CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
@@ F.3 ImageNet-10
Note. CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
@@ F.4 CIFAR-100
Base Model NF C.A. 1% BER 0.5% BER 0.35% BER PA CA (NF) PA (NF) RP PA CA (NF) PA (NF) RP PA CA (NF) PA (NF) RP ResNet18 ConvL 73.7 54.8 11.0 7.7 6.4 65.2 39.0 7.1 18.1 69.4 42.9 6.2 11.4 ConvS 49.7 4.2 2.2 -0.4 70.0 24.5 7.6 3.6 72.1 35.1 7.3 3.7 DeConvL 4.6 55.2 11.9 8.2 7.3 20.9 66.3 38.2 6.9 17.3 31.4 69.2 42.9 5.5 11.4 DeConvS 2.9 32.7 4.0 2.2 -0.6 7.4 68.2 25.9 6.8 5 7.6 71.6 35.8 5.5 4.4 UNetL 50.6 14.5 8.9 10.0 66.2 40.1 6.4 19.2 70.3 46.3 5.5 14.9 UNetS 26.8 4.6 2.5 -0.0 67.1 28.8 6.8 7.9 70.9 38.3 6.4 6.9 ResNet50 ConvL 73.5 63.5 3.2 1.7 0.1 68.4 28.8 6.7 7.6 72.0 40.8 7.5 5.1 ConvS 65.5 3.2 1.6 0.1 71.9 23.1 6.9 1.9 73.0 37.4 8.0 1.7 DeConvL 3.0 59.6 3.2 1.7 0.2 21.3 68.1 28.6 7.0 7.4 35.7 71.7 41.7 7.7 6.1 DeConvS 1.8 61.1 3.2 1.7 0.1 7.0 70.3 25.0 6.7 3.7 8.6 72.8 38.9 7.9 3.3 UNetL 39.0 5.0 1.7 1.9 66.6 36.5 6.2 15.3 70.8 45.3 6.7 9.6 UNetS 47.7 3.4 1.8 0.3 69.1 26.1 6.6 4.8 72.6 39.6 7.8 3.9 VGG11 ConvL 64.8 58.3 19.7 11 11.5 63.1 38.8 9.3 15.0 63.9 42.4 9.0 11.1 ConvS 56.6 10.4 7.4 2.2 62.7 27.9 10 4.0 63.9 41.8 8.3 10.5 DeConvL 8.2 60.3 21.2 11 13.0 23.9 63.9 40.0 9.0 16.2 31.3 64.0 42.8 9.1 11.5 DeConvS 5.7 58.3 11.8 7.9 3.5 9.4 61.9 29.8 9.9 5.9 10 63.5 36.1 10 4.8 UNetL 51.1 22.1 8.2 13.9 61.8 37.8 9.0 13.9 63.5 40.9 9.3 9.6 UNetS 51.9 13.1 7.9 4.9 61.7 29.8 9.7 6.0 63.8 35.7 9.9 4.5 VGG16 ConvL 67.8 51.4 19.2 6.0 12.6 61.8 41.1 5.6 18.7 64.9 44.9 5.3 13.8 ConvS 44.3 6.7 2.3 0.1 63.8 27.5 6.8 5.1 66.0 36.3 6.1 5.1 DeConvL 7.0 53.1 20.8 6.2 14.2 22.4 62.8 42.1 5.5 19.8 31.1 65.0 46.6 5.2 15.5 DeConvS 3.5 23.5 4.8 1.7 -1.8 7.0 62.1 29.9 6.7 7.5 7.2 64.9 38.1 6.3 7.0 UNetL 50.2 25.3 1.7 18.7 61.7 41.3 5.0 18.9 64.8 46.8 4.6 15.7 UNetS 27.7 9.9 2.1 3.3 61.6 31.3 6.3 8.9 65.0 39.8 5.9 8.7 VGG19 ConvL 67.8 59.4 29.2 8.1 18.6 65.6 46.5 6.8 12.5 66.9 49.2 7.4 7.0 ConvS 63.7 14.4 5.1 3.8 66.6 38.3 6.8 4.2 67.7 45.3 8.5 3.2 DeConvL 10.6 60.1 29.6 8.5 19.0 34.0 65.7 46.9 7.1 12.9 42.1 67.3 49.8 7.6 7.6 DeConvS 4.3 60.9 16.1 6.0 5.6 9.6 66.5 39.0 3.7 5.0 9.4 67.7 45.7 8.4 3.6 UNetL 58.7 30.2 8.2 19.6 65.5 46.9 6.5 12.9 67.4 50.0 7.5 7.9 UNetS 59.1 18.0 6.2 7.4 66.3 40.1 8.0 6.1 67.5 46.6 8.4 4.5 • Note. CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
@@ Appendix G Additional Experimental Results on Restricted Access Settings (Transferability)
We find that using VGG19 as a white-box surrogate model has better transferability than ResNet18 for all datasets. In addition, we can observe that some NeuralFuse generators have downward applicability if base models have a similar architecture. In other words, if we try to transfer a generator trained on a large BER (e.g., 1%) to a model with a small BER (e.g., 0.5%), the performance will be better than that of a generator trained with the original BER (e.g., 0.5%). For example, in Table 16, we could find that if we use VGG19 as a source model to train the generator ConvL (1%), the generator could deliver better performance (in terms of PA (NF)) when applied to similar base models (e.g., VGG11, VGG16, or VGG19) under a 0.5% BER, compared to using itself as a source model (shown in Table 12). We conjecture that this is because the generators trained on a larger BER can also cover the error patterns of a smaller BER, and thus they have better generalizability across smaller B.E.Rs.
@@ G.1 CIFAR-10
The results of CIFAR-10 in which NeuralFuse is trained at 1% BER are shown in Table 16.
SM TM BER CA PA ConvL (1%) UNetL (1%) CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 ResNet18 0.5% 92.6 70.1 11.6 89.8 89.5 0.2 19.4 86.6 86.2 0.3 16.1 ResNet50 1% 92.6 26.1 9.4 89.5 36.0 19 9.9 85.2 38.8 19 12.7 0.5% 61.0 10.3 75.1 10 14.1 77.1 5.0 16.1 VGG11 1% 88.4 42.2 11.6 88.4 62.5 8.4 20.3 76.8 61.1 8.5 18.9 0.5% 63.6 9.3 81.0 4.6 17.4 73.7 3.0 10.1 VGG16 1% 90.3 35.7 7.9 89.6 63.3 18 27.6 85.2 59.9 16 24.2 0.5% 66.6 8.1 85.0 3.4 18.4 80.2 4.5 13.6 VGG19 1% 90.5 36.0 12.0 89.6 50.7 22 14.7 85.3 51.1 16 15.1 0.5% 64.2 12.4 80.2 8.7 16.0 76.5 7.8 12.3 VGG19 ResNet18 1% 92.6 38.9 12.4 89.8 61.0 17 22.1 87.0 69.7 11 30.8 0.5% 70.1 11.6 86.1 6.9 16.0 84.2 3.0 14.1 ResNet50 1% 92.6 26.1 9.4 89.9 34.0 19 7.9 87.0 44.2 17 18.1 0.5% 61.0 10.3 76.5 10 15.5 80.7 4.2 19.7 VGG11 1% 88.4 42.2 11.6 89.7 76.5 7.0 34.3 87.1 79.9 5.6 37.7 0.5% 63.6 9.3 88.0 2.1 24.4 85.4 0.8 21.8 VGG16 1% 90.3 35.7 7.9 89.6 75.5 12 39.8 87.2 78.9 7.8 43.2 0.5% 66.6 8.1 88.9 0.6 22.3 86.2 0.3 19.6 VGG19 0.5% 90.5 64.2 12.4 89.8 89.6 8.7 25.4 87.4 86.8 0.4 22.6 • Note. SM: source model, used for training generators; TM: target model, used for testing generators; BER: the bit-error rate of the target model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
@@ G.2 GTSRB
In Tables 17 and 18, we show the results on GTSRB in which NeuralFuse is trained at 1.5% and 1% BER, respectively.
SM TM BER CA PA ConvL (1.5%) UNetL (1.5%) CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 ResNet18 1% 95.5 36.9 16.0 95.7 93.9 1.9 57.0 94.9 94.4 0.4 57.5 0.5% 75.2 12.7 95.7 0.2 20.5 94.8 0.2 19.6 ResNet50 1% 95.0 29.5 16.9 94.4 37.0 22 7.5 94.4 47.1 23 17.6 0.5% 74.0 13.0 77.5 13 3.5 84.8 9.5 10.8 VGG11 1% 91.9 34.9 12.4 92.8 45.2 10 10.3 91.4 50.5 13 15.6 0.5% 64.9 10.8 79.4 5.8 14.5 83.9 4.2 19.0 VGG16 1% 95.2 15.1 6.8 95.4 31.1 13 15.9 94.6 36.8 12 21.7 0.5% 58.8 8.9 84.5 8.3 25.8 86.0 8.6 27.2 VGG19 1% 95.5 36.6 6.8 95.0 56.4 15 19.8 94.3 60.8 15 24.2 0.5% 69.1 11.1 86.9 3.4 17.8 87.7 3.8 18.6 VGG19 ResNet18 1% 95.5 36.9 16.0 88.4 50.3 12 13.4 92.8 63.7 16 26.8 0.5% 75.2 12.7 77.9 7.4 2.7 87.5 3.9 12.3 ResNet50 1% 95.0 29.5 16.9 87.5 29.7 17 0.2 92.5 40.4 21 10.9 0.5% 74.0 13.0 67.9 17 -6.1 77.5 15 3.5 VGG11 1% 91.9 34.9 12.4 89.7 47.1 11 12.2 93.5 60.0 12 25.1 0.5% 64.9 10.8 76.3 5.1 11.4 86.0 3.8 21.1 VGG16 1% 95.2 15.1 6.8 93.0 29.2 15 14.1 93.0 38.5 16 23.4 0.5% 58.8 8.9 75.7 12 16.9 79.9 8.3 21.1 VGG19 1% 95.5 36.6 6.8 95.1 87.4 6.0 50.8 94.6 88.7 5.0 52.1 0.5% 69.1 11.1 92.4 2.4 23.3 92.4 2.2 23.3 • Note. SM: source model, used for training generators; TM: target model, used for testing generators; BER: the bit-error rate of the target model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
SM TM BER CA PA ConvL (1%) UNetL (1%) CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 ResNet18 0.5% 95.5 75.2 12.7 95.7 95.3 0.5 20.1 96.2 95.7 0.3 20.5 ResNet50 1% 95.0 29.5 16.9 94.5 35.6 21 6.1 95.6 42.6 23 13.1 0.5% 74.0 13.0 78.8 13 4.8 87.3 9.0 13.3 VGG11 1% 91.9 34.9 12.4 93.1 45.8 11 10.9 94.0 47.1 14 12.2 0.5% 64.9 10.8 81.8 5.0 16.9 84.2 4.8 19.3 VGG16 1% 95.2 15.1 6.8 95.5 26.5 12 11.4 95.5 32.4 11 17.3 0.5% 58.8 8.9 82.2 9.0 23.4 85.4 6.7 26.6 VGG19 1% 95.5 36.6 6.8 94.9 53.2 14 16.6 95.6 60.9 15 24.3 0.5% 69.1 11.1 85.4 4.5 16.3 87.5 3.7 18.4 VGG19 ResNet18 1% 95.5 36.9 16.0 93.7 53.1 16 16.2 95.0 63.4 18 26.5 0.5% 75.2 12.7 83.9 7.6 8.7 89.7 4.8 14.5 ResNet50 1% 95.0 29.5 16.9 92.8 30.6 18 1.1 95.4 38.9 22 9.4 0.5% 74.0 13.0 74.7 18 0.7 81.5 16 7.5 VGG11 1% 91.9 34.9 12.4 93.7 50.6 11 15.7 95.1 58.9 15 24.0 0.5% 64.9 10.8 82.3 5.1 17.4 87.5 3.7 22.6 VGG16 1% 95.2 15.1 6.8 95.2 27.8 15 12.7 95.2 33.5 14 18.4 0.5% 58.8 8.9 79.0 12 20.2 81.8 7.8 23.0 VGG19 0.5% 95.5 69.1 11.1 96.0 94.0 2.2 24.9 95.4 93.9 2.1 24.8 • Note. SM: source model, used for training generators; TM: target model, used for testing generators; BER: the bit-error rate of the target model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA